OpenStudy (anonymous):

9/25 as a percent how is it done

Parth (parthkohli):

\[\textbf{HINT}\]\[\dfrac{9}{25} = \dfrac{36}{100}\]

OpenStudy (anonymous):

then do you cross multiply

OpenStudy (shubhamsrg):

Yes you cross multiply, tell me what you get?

OpenStudy (anonymous):

900/900

OpenStudy (shubhamsrg):

wow! :O

OpenStudy (shubhamsrg):

Alright, jokes apart, to express a number into percentage, just multiply that by 100.

OpenStudy (anonymous):

324/2500

OpenStudy (shubhamsrg):

what is that ?

Parth (parthkohli):

Hmm... I think it'd be better to multiply \(\dfrac{36}{100}\) by \(100\). Note: NOT \(\dfrac{100}{100}\)

OpenStudy (anonymous):

nothing im so confused and playing jokes on a mentally handicapped person is not funny

OpenStudy (shubhamsrg):

But if another mentally handicapped is cracking those jokes, then it should be ok.

OpenStudy (shubhamsrg):

Nevermind, don't be confused, as I said, or as parth is saying, to convert any number to percentage, just multiply it by 100

OpenStudy (anonymous):

which number

OpenStudy (shubhamsrg):

2